The vertices of a triangle are R (3,0), S (-1,3), and T (0, -2). a) Classify the triangle by side length and b) Find the perimeter of the triangle to the nearest tenth.


1
Expert's answer
2021-11-29T13:20:05-0500


a)


"RS=\sqrt{(-1-3)^2+(3-0)^2}=5"

"RT=\sqrt{(0-3)^2+(-2-0)^2}=\sqrt{13}"

"TS=\sqrt{(-1-0)^2+(-2-3)^2}=\sqrt{26}"

It is a versatile triangle.


b)


"P_{RST}=5+\sqrt{13}+\sqrt{26}\approx13.7"